OpenStudy (anonymous):

Solve: 3m/4-m/3=1/12 A. 5 B.1/5 C.1/6 D.6

OpenStudy (anonymous):

B

jhonyy9 (jhonyy9):

3m/4 -m/3 =1/12 9m-4m=1 5m=1 m=1/5 so B. is right

OpenStudy (anonymous):

Thanks